OpenStudy (isabellemullens):

Which recursive formula can be used to represent the sequence 2, 6, 10, 14, 18,

OpenStudy (jango_in_dtown):

its an AP series with common difference 4 and first term 2. So t_n=a+(n-1)d=2+(n-1)*4=4n-2

OpenStudy (loser66):

*+ 4n+2